OpenStudy (anonymous):

Check for convergence or divergence of the integral (x+1)/(x^2) dx from -1 to 1

OpenStudy (anonymous):

\[\int\limits_{-1}^{1} \frac{ (x+1) }{ x^2 } dx\]

OpenStudy (anonymous):

\[\int \frac{1}{x}+\frac{1}{x^2}dx\] not likely to converge

OpenStudy (anonymous):

since the anti derivative of the first part is the log, and the log of zero gives minus infinity